OpenStudy (anonymous):

Let f(x) = x + 7 and g(x) = x – 2. What is f(x) • g(x)? x2 – 5x + 14 x2 + 5x – 14 x2 + 5x – 5 x2 + 9x + 5

OpenStudy (anonymous):

Multiply 'em up, guuurl. Remember to FOIL.

OpenStudy (anonymous):

\[f(x).g(x) = (x+7)(x-2) = x(x-2) + 7(x-2)\] Distribute x and 7 to the corresponding parenthesis and then solve..

OpenStudy (anonymous):

use foil :)

OpenStudy (anonymous):

(x+7)(x-2) x(x-2)+7(x-2) x^2-2x+7x-14 x^2+5x-14

OpenStudy (anonymous):

didnt get?????????

OpenStudy (anonymous):

Kaafi moti ho gyi ho swati...

OpenStudy (anonymous):

haaaaaaaaaa kya????

OpenStudy (anonymous):

English me kahunga to sab jaan jaaenge.. Ha ha ha..

OpenStudy (anonymous):

shut up

OpenStudy (anonymous):

Gussa maan gyi kya?? Arey main to majaak kar raha tha bas..